OpenStudy (anonymous):

(sqrt(2)+2(sqrt(-32)))-(sqrt(8))-(sqrt(8)-(sqrt(-18))) Simplify the answers i can choice from are -sqrt(2)+11i(sqrt(2)), sqrt(2)-3isqrt(2), 3sqrt(2)-8isqrt(2),-2sqrt(2)+5isqrt(2)

OpenStudy (anonymous):

you sure its 2(sqrt(-32)

OpenStudy (anonymous):

could it be 32 not -32?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its -32

OpenStudy (anonymous):

\[2 \sqrt{-32}-\sqrt{-18}+\sqrt{2}-\sqrt{8}-\sqrt{-18}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes or no?

OpenStudy (anonymous):

no its in the order order of how i wrote it

OpenStudy (anonymous):

\[(\sqrt{2}+2\sqrt{-32})-(\sqrt{8}-\sqrt{-18})\]

OpenStudy (anonymous):

\[(\sqrt{2}+2\sqrt{-32})-(\sqrt{8}-\sqrt{-18})\]

OpenStudy (anonymous):

\[\left(\sqrt{2}+2\sqrt{-32}\right)-\left(\sqrt{8}-\sqrt{-18}\right)\] \[\left(\sqrt{2}+2\sqrt{-2}\sqrt{16}\right)-\left(\sqrt{8}-\sqrt{9}\sqrt{-2}\right)\] \[\left(\sqrt{2}+8\sqrt{-2}\right)-\left(\sqrt{8}-3\sqrt{-2}\right)\]

OpenStudy (anonymous):

\[\left(\sqrt{2}+8\sqrt{2}\sqrt{-1}\right)-\left(\sqrt{8}-3\sqrt{2}\sqrt{-1}\right)\]

OpenStudy (anonymous):

\[\left.\sqrt{2}+8\sqrt{2}i\right)-\left(\sqrt{8}-3\sqrt{2}i\right)\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

you can simplify that further

OpenStudy (anonymous):

could you please simplify it all the way?

OpenStudy (anonymous):

\[\left(8 \sqrt{2} i+\sqrt{2}\right)-\left(2 \sqrt{2}-3 \sqrt{2} i\right)\]